OpenStudy (anonymous):

how do you solve y = 3/ sqrt(16-x^2) ?

OpenStudy (anonymous):

\[y=\frac{3}{\sqrt{16+x^2}}\] what are you solving for?

OpenStudy (anonymous):

The area of the region between 0 and 2.

OpenStudy (anonymous):

you are looking for \[\int^2_0 \frac{3}{\sqrt{16-x^2}}dx \]?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I know that it has to do with the sine identity, but I keep getting the wrong answer using that. My answer is never an angle that is correct

OpenStudy (amistre64):

x = 4sin(t)

OpenStudy (amistre64):

dx = 4cos(t) dt

OpenStudy (anonymous):

antiderivative of \[\frac{1}{\sqrt{a^2-x^2}}\] is \[sin^{-1}(\frac{x}{a})\]

OpenStudy (amistre64):

3*4 cos(t) [S] --------- dt right? 4 cos(t)

OpenStudy (anonymous):

\[1/a *\arcsin(x/a) ?\]

OpenStudy (amistre64):

[S] 3 dt -> 3t ... what does t revert back to tho? x = 4sin(t) x/4 = sin(t) sin^-1(x/4) = t if i did it right :)

OpenStudy (amistre64):

3 * sin^-1(x/4)

OpenStudy (anonymous):

I pulled out the three and had (1/4)arcsin(x/4) evaluated from 0 to 2

OpenStudy (anonymous):

evaluate at upper limit, get \[sin^{-1}(\frac{2}{4})=sin^{-1}(\frac{1}{2})=\frac{\pi}{6}\] plug in 0 get 0 so answer is \[\frac{\pi}{6}\]

OpenStudy (anonymous):

Did you multiply by the 3?

OpenStudy (anonymous):

oops forgot about the ''3" answer is \[\frac{\pi}{2}\] sorry

OpenStudy (anonymous):

So when I get a decimal, how can I revert it make to degrees/radians so that the computer will except it?

OpenStudy (amistre64):

\[frac{test}{ing}\]

OpenStudy (amistre64):

lol... that didnt work out like i thought

OpenStudy (anonymous):

what's that?

OpenStudy (amistre64):

pi 180 -- * ----- = 90 degrees... 2 pi

OpenStudy (anonymous):

pi/2 is right

OpenStudy (anonymous):

the area is the area. As a function this is a function of numbers, not degrees or radians. It corresponds to the radian measure of the angle, not the degree measure. Always work in 'radians' if you are thinking of angles. But they are just numbers.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Thanks everyone for your help!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

I hope this is clear. the area is \[\frac{\pi}{2}\] square units.

OpenStudy (anonymous):

Yes, and your answer was right.